### v2

# Nuclear state and level densities of actinides in the shell-model Monte Carlo

Abstract
Actinides are of great interest in astrophysics and technology applications since they can fission. However, the microscopic calculation of their statistical properties in the presence of correlations poses a major theoretical challenge. The configuration-interaction shell-model is a suitable framework to calculate these properties but the required large model spaces are beyond the reach of conventional diagonalization methods. The shell-model Monte Carlo (SMMC) method enables calculations in very large model spaces and was applied to nuclei as heavy as the lanthanides. Here, we extend the SMMC method to the actinides. Fifteen even-even and odd-mass actinides 232Th, $^{\textrm{234-239}}$U, $^{\textrm{240-243}}$Pu, $^{\textrm{246-248}}$Cm, and 250Cf are studied using a single-particle model space that is larger than one major shell each for protons and neutrons, with a total dimension of the many-particle space as large as $10^{32}$. We calculate nuclear state densities of these actinides and find they are strongly enhanced in comparison with mean-field densities. We use spin projection methods to calculate nuclear level densities and average $s$-wave neutron resonance spacings, both of which are found to be in good agreement with experiments.
Introduction.— The nuclear level density (NLD) is a fundamental statistical property of nuclei with importance to compound-nucleus reactions [1], astrophysical abundances and processes [2], and more. Theoretical calculations of the statistical properties of heavy open-shell nuclei have required either highly truncated model spaces, or the use of mean-field approximations (e.g., density function theory [3]) which miss important correlations between nucleons. Accurate calculations of statistical properties in the actinide mass region ($A\sim 240$) require configuration-interaction (CI) shell model spaces larger than $10^{30}$, while computational costs currently limit conventional diagonalization methods to model spaces around $10^{11}$ [4, 5]. The shell-model Monte Carlo (SMMC) [6] method enables microscopic calculations of nuclear properties in the presence of correlations in shell model spaces that are many orders of magnitude larger than what can be treated in conventional methods. Previously, the SMMC method has been applied to nuclei as heavy as the lanthanides with $A\sim 160$ [7, 8, 9, 10, 11, 12, 13].
Several experimental methods have been used to measure NLDs. The Oslo method [14, 15, 16] has been used to measure the NLDs of nuclei across a broad region of the nuclear chart at excitation energies below their neutron separation energies $S_{n}$. In particular, the Oslo method has been used to measure NLDs in several actinides [17, 18, 19, 20].
Actinides can fission and the deformation dependence of their NLDs is an important input to models of nuclear fission [21, 22]. Additionally, the quadrupole deformation of 238U, the largest nucleus collided at the Relativistic Heavy Ion Collider, can impact observables like the transverse momentum distributions of particles detected in relativistic heavy ion collisions [23, 24, 25, 26].
In this work we present the first applications of the SMMC method to actinides. The larger model space and necessity to calculate observables at very large inverse temperatures $\beta=1/T$ (in order to cool the nucleus to its ground state) present additional technical challenges compared with previous applications of SMMC in lanthanide nuclei. The calculated NLDs are found to be in similarly good agreement with experiment as has been seen in the lighter lanthanides. This work constitutes the first step towards studying statistical properties of actinides that are much beyond the reach of conventional large-scale shell-model calculations.
Shell-model Monte Carlo.— The SMMC method is based on the canonical ensemble describing a nucleus at finite temperature $T$. The Gibbs operator $e^{-\beta\hat{H}}$ (where $\beta=1/T$) is transformed by the Hubbard-Stratonovich (HS) transformation [27, 28] into a functional integral of one-body propagators $\hat{U}_{\sigma}$ that describe non-interacting nucleons in a set of external, imaginary-time-dependent fields $\sigma(\tau)$ with $0\leq\tau\leq\beta$
$e^{-\beta\hat{H}}=\int D[\sigma]G_{\sigma}\hat{U}_{\sigma},$ (1)
where $G_{\sigma}$ is a Gaussian weight and $D[\sigma]$ is the integral measure.
The thermal average of an observable $\hat{O}$ at inverse temperature $\beta$ is then given by
$\langle\hat{O}\rangle=\frac{\textrm{Tr}(\hat{O}e^{-\beta\hat{H}})}{\textrm{Tr}(e^{-\beta\hat{H}})}=\frac{\int D[\sigma]G_{\sigma}\langle\hat{O}\rangle_{\sigma}\textrm{Tr}\hat{U}_{\sigma}}{\int D[\sigma]G_{\sigma}\textrm{Tr}\hat{U}_{\sigma}}\;,$ (2)
where $\langle\hat{O}\rangle_{\sigma}=\textrm{Tr}(\hat{O}\hat{U}_{\sigma})/\textrm{Tr}\,\hat{U}_{\sigma}$. The integration over the auxiliary fields is evaluated using Monte Carlo methods, in which the fields $\sigma(\tau)$ are samples with the positive-definite weight $W_{\sigma}\equiv G_{\sigma}|\textrm{Tr}\hat{U}_{\sigma}|$. In this work we use good-sign interactions, for which the Monte-Carlo sign $\Phi_{\sigma}\equiv\textrm{Tr}\hat{U}_{\sigma}/|\textrm{Tr}\hat{U}_{\sigma}|$ is always positive in the framework of the grand-canonical ensemble.
Quantities in the integrand of Eq. (2) can be calculated using the matrix representation ${\bf U}_{\sigma}$ of the propagator $\hat{U}_{\sigma}$ in the single-particle model space, reducing the required matrix algebra to matrices the size of the single-particle space ($\sim 100$) rather than the combinatorially-large spaces required in direct diagonalization methods. For example, the grand-canonical trace of $\hat{U}_{\sigma}$ is given by
$\textrm{Tr}\,\hat{U}_{\sigma}=\det(\boldsymbol{1}+{\bf U}_{\sigma}).$ (3)
The imaginary-time interval $0\leq\tau\leq\beta$ is discretized into $N_{\tau}$ time slices of width $\Delta\beta=\beta/N_{\tau}$. For any given value of $\beta$, thermal averages of observables are calculated for $\Delta\beta=1/32$ MeV-1 and $\Delta\beta=1/64$ MeV-1. The continuous time limit $\Delta\beta\rightarrow 0$ of the expectation value of an observable is then determined by linear extrapolation of the expectation values computed at these two values of $\Delta\beta$.
We work in the canonical ensemble with fixed proton and neutron numbers so the grand-canonical traces in Eq. (2) are replaced by canonical traces at fixed particle number $A$, i.e., we substitute $\textrm{Tr}\rightarrow\textrm{Tr}_{A}$. This is accomplished by an exact particle-number projection onto $A$ valence nucleons via a discrete Fourier transform [29, 30]
$\textrm{Tr}_{A}\hat{U}_{\sigma}=\frac{e^{-\beta\mu A}}{N_{s}}\sum_{m=1}^{N_{s}}e^{-i\varphi_{m}A}\det(\boldsymbol{1}+e^{i\varphi_{m}+\beta\mu}\bf{U_{\sigma}})\;,$ (4)
where $N_{s}$ is the number of single-particle states, $\varphi_{m}=2\pi m/N_{s}$ are quadrature points and $\mu$ is a real chemical potential introduced to stabilize the Fourier sum. In practice, we use two particle-number projections onto fixed proton and neutron numbers.
For good-sign interactions (in the grand-canonical ensemble), the projection onto an even number of particles keeps the good sign, but the projection on an odd number of particles leads to a sign problem at low temperatures. This makes direct calculations of the ground-state energy of odd-mass nuclei impossible. Here we use the partition function extrapolation method introduced in Ref. [31] to calculate the ground-state energy of the odd-mass actinides (see the Supplemental Material [32]).
Model space and interaction for actinides.— A CI shell-model Hamiltonian has been developed to describe statistical properties of actinides. The model space consists of spherical orbitals and single-particle energies computed from a central Woods-Saxon potential with spin-orbit coupling. The doubly-magic 208Pb is used as the inert core. For protons, the valence space consists of the 82-126 shell plus the $1g_{9/2}$ orbital, while the neutron valence space consists of the 126-184 shell plus the $1h_{11/2}$ orbital [33]. For the heaviest nuclei studied in this work, the total m-scheme model space dimension is greater than $10^{32}$.
The interaction consists of a pairing-plus-multipole interaction which has a good Monte-Carlo sign and includes the dominant collective components of realistic nuclear interactions [34]. It is given by
$\sum_{\nu}g_{\nu}\hat{P}_{\nu}^{\dagger}\hat{P}_{\nu}-\sum_{\lambda}\chi_{\lambda}:(\hat{\mathcal{O}}_{\lambda p}+\hat{\mathcal{O}}_{\lambda n})\cdot(\hat{\mathcal{O}}_{\lambda p}+\hat{\mathcal{O}}_{\lambda n}):,$ (5)
where $\nu=p,n$ is the nucleon species, $\lambda=2,3,4$ corresponds to the $2^{\lambda}$-pole interactions, $P_{\nu}^{\dagger}$ is the monopole pair creation operator, $\hat{\mathcal{O}}_{\lambda\nu}$ are the surface-peaked multipole operators, and :::: denotes normal ordering. The multipole strengths are defined as $\chi_{\lambda}=\chi k_{\lambda}$ where $\chi$ is determined self-consistently [35] and $k_{\lambda}$ are renormalization factors that take into account the effects of the inert core. The interaction coefficients are parameterized by the following $N$-dependent functional forms:
[…]
These coefficients were determined such that they reproduce known odd-even mass staggerings and models of nuclear state densities fitted to experimental data. See the Supplemental Material [32] and references therein [36, 37, 38, 39, 40, 41] for more details.
State and level densities of actinides.— The nuclear state density (NSD) at energy $E$ is the inverse Laplace transform of the nuclear partition function $Z(\beta)$. This transform is numerically ill-behaved, and instead we compute it in the saddle-point approximation where it provides the average NSD
$\rho(E)=\frac{1}{2\pi i}\int_{-i\infty}^{+i\infty}d\beta e^{\beta E}Z(\beta)\approx\sqrt{\frac{\beta^{2}}{2\pi C}}e^{S(E)}\;.$ (7)
In Eq. (7) $S$ and $C$ are, respectively, the canonical entropy and heat capacity. The logarithm of the partition function $\ln Z(\beta)$ is computed by integrating the thermodynamic relation $\partial\ln Z/\partial\beta=E(\beta)\equiv\langle H\rangle$ and $\beta$ is determined as a function of $E$ using the saddle-point condition $E(\beta)=E$. The canonical entropy is then given by $S(E)=\ln Z+\beta E$, and the canonical heat capacity $C$ is computed with the method of Ref. [42] using correlated errors to reduce statistical errors. Thermal energies $E(\beta)$ are calculated up to $\beta=40$ MeV-1 for even-even nuclei and $\beta=8$ MeV-1 for odd-mass nuclei. The corresponding temperature for the even-even nuclei is sufficiently low so that the ground-state energy $E_{0}$ can be determined by fitting the thermal energy to a form based on a ground-state rotational band. Thermal energies $E(\beta)$ are calculated up to $\beta=40$ MeV-1 for even-even nuclei and $\beta=8$ MeV-1 for odd-mass nuclei. The corresponding temperature for the odd-mass nuclei is not sufficiently low to directly extract $E_{0}$ and we use the recently developed partition function extrapolation method [31]. The excitation energy is then calculated using $E_{x}=E-E_{0}$.
Figure 1: Comparison of the NSDs for actinides calculated in the SMMC (solid circles), HFB (solid lines) and the experimentally determined [43] BBF (8) (dashed lines). The red triangles are the NSDs at the neutron separation energies $S_{n}$ and the histograms describe experimentally known states with assigned spins up to 1.2 MeV [44].
In comparing the SMMC NSDs with experimental data, we used the back-shifted Bethe formula (BBF) [45]
$\rho_{\textrm{BBF}}(E_{x})=\frac{\sqrt{\pi}e^{2\sqrt{a(E_{x}-\Delta)}}}{12a^{1/4}(E_{x}-\Delta)^{5/4}},$ (8)
where $a$ and $\Delta$ are parameters determined by a fit to low-lying state counting and neutron resonance average spacing data [43] taken from RIPL-3 [44]. Figure 1 compares the NSDs for the actinides calculated with the SMMC (solid circles) and in the finite-temperature Hartree-Fock-Boguliubov (HFB) approximation [46] (solid lines) using the code HF-SHELL [^1] HF-SHELL only supports calculations in even-even nuclei. [47]. While the SMMC shows good agreement with the experimental results, the HFB results are lower by about an order of magnitude. The enhancement of the SMMC NSD relative to the HFB density is attributed to rotational states which are missed by the HFB as these nuclei are strongly deformed ($0.24<\beta_{2}<0.3$). Previous results have shown that the SMMC and HFB show better agreement with each other as the nuclear deformation decreases [11].
Figure 2: The enhancement factors $K=\rho_{\textrm{SMMC}}/\rho_{\textrm{HFB}}$ of 234U, 236U, and 238U versus excitation energy $E_{x}$. The dashed line corresponds to $K=1$.
The contribution of these rotational bands is quantified by the enhancement factor $K=\rho_{\textrm{SMMC}}/\rho_{\textrm{HFB}}$. The energy dependence of $K$, shown in Fig. 2 for the even-even uranium isotopes, displays some important features. For the excitation energies of interest in this work, the enhancement factors are between 10 and 25, indicating a large number of rotational bands that are missed in the HFB (the latter takes into account only the intrinsic band head states). The rapid decrease in $K$ starting around $E_{x}\sim 15$ MeV is due to the transition from deformed to spherical shapes of the nuclei. At the highest excitation energies shown in Fig. 2, when the nuclei are essentially spherical, $K\approx 1$ indicating agreement between the HFB and the SMMC as the rotational bands have disappeared. The large enhancement at very low excitation energies is an unphysical artifact of the non-conservation of particle number in the HFB.
Figure 3: Comparison of the SMMC NLDs (solid circles) with the Oslo method experiments [17, 19, 20] (solid squares) for the same actinides as in Fig. 1. The red triangles are the NLDs at the neutron separation energies $S_{n}$ and the histograms are experimentally known levels up to 1.2 MeV [44].
Nucleus $S_{n}$ (MeV) $J_{t}$ $\tilde{\rho}(S_{n})$ ($10^{6}$ MeV-1) $\sigma(S_{n})$ $D_{0}$ (eV)
  Exp.     SMMC rigid body    SMMC   Exp.     SMMC
232Th 6.438 5/2 30. $\pm$ 8.222The experimental level density for 232Th was estimated from systematics in Ref. [17]. 20.5 $\pm$ 3.2 rigidsrigids 8.32 $\pm$ 0.19 1.21 $\pm$ 0.20
237U 5.126 0 9.3 $\pm$ 1.9 5.53 $\pm$ 2.32 7.95 8.05 $\pm$ 0.21 14.0 $\pm$ 1.0 23.6 $\pm$ 10.0
238U 6.154 1/2 20. $\pm$ 6. 19.7 $\pm$ 2.8 8.11 8.39 $\pm$ 0.17 3.5 $\pm$ 0.8 3.62 $\pm$ 0.53
239U 4.806 0 6.1 $\pm$ 1.2 3.19 $\pm$ 0.69 7.83 7.93 $\pm$ 0.22 20.3 $\pm$ 0.6 39.7 $\pm$ 8.9
240Pu 6.534 1/2 32.7 $\pm$ 6.6 24.7 $\pm$ 3.5 8.32 8.58 $\pm$ 0.17 2.20 $\pm$ 0.09 3.01 $\pm$ 0.44
243Pu 5.034 0 7.87 $\pm$ 1.93 3.54 $\pm$ 2.11 7.96 8.07 $\pm$ 0.22 13.5 $\pm$ 1.5 37.1 $\pm$ 22.2
Table 1: The SMMC NLDs $\bar{\rho}(S_{n})$ at the neutron separation energies $S_{n}$ are compared with the respective NLDs determined from neutron resonance spacing data [17, 19, 20]. We also compare the average s-wave neutron resonance spacing $D_{0}$ calculated with the SMMC (12) with the experimentally measured values [44].
The NSD counts the total number of states including the $2J+1$ magnetic degeneracy for each level of spin $J$. However, the NLD, measured in experiments such as the Oslo method, counts each level with spin $J$ only once. The NLD can be computed directly through the use of spin projection [48]. For each level of spin $J$ in an even-mass (odd-mass) nucleus there exists exactly one state with $M=0$ ($M=1/2$). Thus by projecting thermal observables onto $M=0$ ($M=1/2$) for even-mass (odd-mass) nuclei the contribution of only one state from each level is counted [49].
The trace over an observable $\hat{O}$ at fixed spin component $M$ is calculated using a discrete Fourier transform
[…]
where $\varphi_{k}=\frac{k\pi}{J_{s}+1/2}$ are quadrature points with $k=-J_{s},...,J_{s}$ and $J_{s}$ being the maximum possible spin in the model space. The projected densities $\rho_{M}$ can be calculated using an equation similar to Eq. (7) but with the projected thermal energies $E_{M}$ replacing the thermal energy $E$. The NLDs $\tilde{\rho}$ are then given by $\tilde{\rho}=\rho_{M=0}$ ($\tilde{\rho}=\rho_{M=1/2}$) for even-mass (odd-mass) nuclei [49]. The Fourier sum in Eq. (9) can lead to non-positive $M$-projected traces for non-zero $M$, resulting in a sign problem for $M\neq 0$ at very low temperatures. Figure 3 shows the NLDs calculated with the SMMC. The SMMC results show good agreement with Oslo method data (solid squares), counting of low-lying levels (red histograms), and neutron resonance data (red triangle) at the neutron separation energy $S_{n}$. Table 1 provides the calculated and measured values of the NLDs at the neutron separation energies. The NLDs at these energies are needed explicitly by the Oslo method, and are usually estimated via neutron resonance spacing data and the spin-cutoff model.
The spin (as well as parity) distribution of level densities is another important input in the calculation of nuclear reaction rates [50]. This spin distribution is also necessary for calculating total level densities from the measured average neutron resonance spacings [14]. Typically, it is assumed that at the neutron resonance energy the spin distribution follows the spin-cutoff model [51]
$\frac{\rho_{J}(E_{x})}{\rho(E_{x})}=\frac{2J+1}{2\sqrt{2\pi}\sigma^{3}}e^{-J(J+1)/2\sigma^{2}},$ (10)
where $\sigma$ is the spin-cutoff parameter, which depends on $E_{x}$ and is related to a thermal moment of inertia $I$ by $\sigma^{2}=IT/\hbar^{2}$.
Assuming that both parities have equal densities at $S_{n}$, the average s-wave neutron resonance spacing $D_{0}$ of the target nucleus in the neutron resonance experiment (i.e., the nucleus with one neutron fewer than the nucleus whose NLD is being considered) can be related to the spin-dependent level densities by
$\frac{2}{D_{0}}=\begin{cases}{\rho}_{1/2}(S_{n})&\textrm{if }J_{t}=0\\ ({\rho}_{J_{t}-1/2}(S_{n})+{\rho}_{J_{t}+1/2}(S_{n}))&\textrm{if }J_{t}\neq 0,\end{cases}$ (11)
where $J_{t}$ is the ground-state spin of the target nucleus. Further assuming that the spin distribution is well described by the spin-cutoff model (10), we find
$\tilde{\rho}(S_{n})=\frac{2\sigma^{2}}{D_{0}((J_{t}+1)e^{-(J_{t}+1)^{2}/2\sigma^{2}}+J_{t}e^{-J_{t}^{2}/2\sigma^{2}})}.$ (12)
In the SMMC, the spin-projected level densities are calculated from the $M$-projected level densities using $\rho_{J}=\rho_{M=J}-\rho_{M=J+1}$. By computing the spin-projected NLDs, we can determine the spin-distribution $\rho_{J}/\rho$ at fixed excitation energy $E_{x}$. We find that at excitation energies near $S_{n}$, the spin distribution is very well described by the spin cutoff model. A fit is then performed to determine the spin-cutoff parameter $\sigma(S_{n})$ with high accuracy and used to calculate the average neutron resonance $s$-wave spacing $D_{0}$.
The fitted values of $\sigma(S_{n})$ shown in Table 1 are found to be within a few percent of their rigid-body values, as well as from the empirical model of Refs. [52, 53] that was used in the Oslo method analyses. For the even-even nuclei, we find the SMMC results for $\tilde{\rho}(S_{n})$ and $D_{0}$ to agree with their experimental values. For the odd-mass nuclei, the SMMC values of $\tilde{\rho}(S_{n})$ and $D_{0}$ are, respectively, lower and higher than their corresponding experimental values. We note however that the uncertainties in these quantities for the odd-mass nuclei are much larger due to the larger uncertainties in the ground-state energies.
Conclusion.— In conclusion, we have carried out the first microscopic calculations of NSDs and NLDs in the actinide region using the SMMC. The SMMC is the only shell-model method currently able to handle the large model spaces required for this mass region. Our results are in agreement with experimental level counting data at low excitations, neutron resonance data, and recent Oslo method results. The SMMC spin projection method enables predictions of the spin cutoff parameter and neutron resonance average spacing in nuclei that have not yet been studied experimentally.
Acknowledgments.— This work was supported in part by the U.S. DOE grant No. DE-SC0019521. Computational resources used for this work were provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. DOE Office of Science User Facility operated under Contract No. DE-AC02-05CH11231.
The data that support the findings of this article are openly available [54].
Appendix A End matter
Results for other actinides.— In the following we present results for nine other actinides for which Oslo method experiments have not been performed but for which neutron resonance data and low-lying level counting data exist. The NSDs, NLDs, and spin distributions at the respective neutron separation energies have been calculated for these nuclei using the same methods as those presented in the main text.
Figures 4 and 5 show the NSDs and NLDs, respectively, for these nuclei. The NSDs of these nuclei show similarly good agreement with the BBF densities as the actinides in the main text. Table 2 provides the corresponding experimental and calculated values of the NLDs, spin-cutoff parameters and $D_{0}$ at the neutron separation energies.
Figure 4: As in Fig. 1, but for other actinides.
Figure 5: As in Fig. 3, but for other actinides.
Nucleus $S_{n}$ (MeV) $J_{t}$ $\tilde{\rho}(S_{n})$ ($10^{6}$ MeV-1) $\sigma(S_{n})$ $D_{0}$ (eV)
  Exp.     SMMC rigid body    SMMC   Exp.     SMMC
234U 6.845 5/2 46.8 $\pm$ 8.9 49.4 $\pm$ 5.6 rigid 8.24 rigid 8.64 $\pm$ 0.13 0.52 $\pm$ 0.02 0.54 $\pm$ 0.11
235U 5.298 0 11.7 $\pm$ 2.5 6.74 $\pm$ 0.74 8.07 8.11 $\pm$ 0.15 $11.2\pm 0.8$ 19.7 $\pm$ 2.3
236U 6.546 7/2 41.9 $\pm$ 7.8 38.3 $\pm$ 3.2 8.15 8.56 $\pm$ 0.12 0.45 $\pm$ 0.03 0.53 $\pm$ 0.07
241Pu 5.242 0 $10.2\pm 2.1$ 4.48 $\pm$ 0.57 8.12 8.12 $\pm$ 0.18 $13.0\pm 0.5$ 19.7 $\pm$ 4.0
242Pu 6.310 5/2 26.3 $\pm$ 5.4 25.9 $\pm$ 5.3 8.27 8.61 $\pm$ 0.17 0.93 $\pm$ 0.08 1.02 $\pm$ 0.23
246Cm 6.458 7/2 24.6 $\pm$ 4.6 26.7 $\pm$ 4.3 8.30 8.60 $\pm$ 0.18 0.79 $\pm$ 0.05 0.78 $\pm$ 0.18
247Cm 5.156 0 $4.71\pm 1.22$ 3.29 $\pm$ 0.42 8.37 8.02 $\pm$ 0.20 30.0 $\pm$ 5.0 39.4 $\pm$ 5.4
248Cm 6.213 9/2 13.9 $\pm$ 4.1 19.8 $\pm$ 2.4 8.32 8.47 $\pm$ 0.18 1.2 $\pm$ 0.3 0.87 $\pm$ 0.11
250Cf 6.624 9/2 22.7 $\pm$ 4.8 41.1 $\pm$ 4.9 8.51 8.66 $\pm$ 0.16 0.76 $\pm$ 0.10 0.43 $\pm$ 0.06
Table 2: As in Table 1, but for other actinides